# Finite almost simple groups whose holomorph contains a solvable regular subgroup

### 25 Dec 2023 v1

@@ Abstract.
In our previous paper, we gave a complete list of the finite non-abelian simple groups whose holomorph contains a solvable regular subgroup. In this paper, we refine our previous work by considering all finite almost simple groups. In particular, our result yields a complete characterization of the finite almost simple groups which occur as the type of a Hopf-Galois structure on a solvable extension, or equivalently, the additive group of a skew brace with solvable multiplicative group.
@@ 1. Introduction
Let $G$ and $N$ be two finite groups of the same order. We say that the pair $(G,N)$ is realizable if the holomorph of $N$ contains a regular subgroup which is isomorphic to $G$. Recall that the holomorph of $N$ is defined to be
@@
For a pair $(G,N)$ to be realizable, it is conjectured that if $G$ is insolvable then $N$ is also insolvable (see [8, 3] for some progress on this). On the other hand, the converse is known to be false – there exist realizable pairs $(G,N)$ for which $G$ is solvable but $N$ is insolvable. It seems to be a natural problem to try and determine the insolvable groups $N$ that can occur in such a pair. In a previous paper [9, Theorems 1.2 & 1.3], we proved that:
@@ 1. $(a)$
    such that $P=AB$ is factorized by some solvable subgroups $A,\,B$ which satisfy $A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$.
@@ 2. $(b)$
    such that $P=AB$ is factorized by some solvable subgroups $A,\,B$ which satisfy $A\cap B=1,\,A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$, and $A$ splits over $A\cap\mathrm{Inn}(N)$, then $(G,N)$ is realizable for some solvable group $G$.
@@ 6. $(6)$
$\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{(2,q-1)}\times C_{f}.$
@@
Therefore, the converse of Theorem 1.1 holds with only five exceptions of $N$ up to isomorphism. The exceptions are the following:
@@
2. Proof for $\mathrm{PSL}_{3}(3),\,\mathrm{PSL}_{3}(4),\,\mathrm{PSL}_{3}(8),\,\mathrm{PSU}_{3}(8),\,\mathrm{PSU}_{4}(2),\,\mathrm{M}_{11}$
@@
  //A list to contain the P which satisfies the conditions
@@
This proves cases (1), (3), (4), (5), (6) of Theorem 1.3.
3. Proof for $\mathrm{PSL}_{2}(q)$ with $q\neq 2,3$ a prime power
@@
We may regard $\mathrm{PSL}_{2}(q)$ as a normal subgroup of $\mathrm{PGL}_{2}(q)$ via the natural embedding, and in particular $\mathrm{PGL}_{2}(q)$ as a subgroup of $\mathrm{Aut}(\mathrm{PSL}_{2}(q))$ which acts on $\mathrm{PSL}_{2}(q)$ via conjugation. It is well-known that
@@
Clearly $N$ is contained $P$. For the choices of $A,\,B$, we separate into cases.
@@
  • $q\equiv 3\pmod{4}:$ Here $A=D\rtimes E$. Since $q=p^{f}$, the exponent $f$ must be odd in this case, and $\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{2}\times C_{f}\simeq C_{2f}$ is cyclic. We then see that there are only two possibilities for $N$, namely

### 21 Mar 2024 v2

@@ Abstract.
In our previous paper, we gave a complete list of the finite non-abelian simple groups whose holomorph contains a solvable regular subgroup. In this paper, we refine our previous work by considering all finite almost simple groups. In particular, our result yields a complete characterization of the finite almost simple groups which occur as the type of a Hopf-Galois structure on a solvable extension, or equivalently, the additive group of a skew brace having a solvable multiplicative group.
@@ 1. Introduction
Let $G$ and $N$ be two finite groups of the same order. We say that the pair $(G,N)$ is realizable if the holomorph of $N$ contains a regular subgroup that is isomorphic to $G$. Recall that the holomorph of $N$ is defined to be
@@
For the pair $(G,N)$ to be realizable, it is conjectured that if $G$ is insolvable then $N$ is also insolvable (see [8, 3] for some progress on this). On the other hand, the converse is known to be false – there exist realizable pairs $(G,N)$ for which $G$ is solvable but $N$ is insolvable. It seems to be a natural problem to try and determine the insolvable groups $N$ that can occur in such a pair. In a previous paper [9, Theorems 1.2 & 1.3], we proved that:
@@ 1. $(a)$
    such that $P=AB$ is factorized by some solvable subgroups $A$ and $B$ that satisfy the equality $A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$.
@@ 2. $(b)$
    such that $P=AB$ is factorized by some solvable subgroups $A$ and $B$ that satisfy $A\cap B=1,\,A\mathrm{Inn}(N)=B\mathrm{Inn}(N)$, and $A$ splits over $A\cap\mathrm{Inn}(N)$, then $(G,N)$ is realizable for some solvable group $G$.
@@ 6. $(6)$
$\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{\gcd(2,q-1)}\times C_{f}.$
@@
Therefore, the converse of Theorem 1.1 holds with only five exceptions of $N$, as follows. Here, only the isomorphism class of $N$ matters, so it suffices to consider the subgroups of $\mathrm{Out}(\mathrm{Soc}(N))$ up to conjugacy.
@@
2. Proof of cases (2) to (6)
@@
  //A list to contain the P that satisfies the conditions
@@
3. Proof of case (1)
@@
We may regard $\mathrm{PSL}_{2}(q)$ as a normal subgroup of $\mathrm{PGL}_{2}(q)$ via the natural embedding, and in particular $\mathrm{PGL}_{2}(q)$ as a subgroup of $\mathrm{Aut}(\mathrm{PSL}_{2}(q))$ that acts on $\mathrm{PSL}_{2}(q)$ via conjugation. It is well-known that
@@
Clearly $P$ contains $N$. For the choices of $A,\,B$, we separate into cases.
@@
  • $q\equiv 3\pmod{4}:$ Here $A=D\rtimes E$. Since $q=p^{f}$, the exponent $f$ must be odd in this case, and so
    $\mathrm{Out}(\mathrm{PSL}_{2}(q))\simeq C_{2}\times C_{f}\simeq C_{2f}$
    is cyclic. We then see that there are only two possibilities for $N$, namely